OpenStudy (anonymous):

help with SAT math please jason buys and sells baseball cards to earn extra amoney. he purchased the card of a player who performed so well throughtout the year that jason was able to mark the card up 20% more than what he orginally paid. when he sold the card, he gave one of his good friends a discount of 10% off his already marked up price. what was Jason's percent profit on the card?

OpenStudy (anonymous):

let Jason's buying price be 100 20% markup = 20% of 100 = 20 so selling price/marked price = 120 discount if 10% on marked price discount = 10% 0f 120 = 12 so final selling price is 120 - 12 = 108 so profit = 108-100 = 8 profit percent = 8/100 * 100% = 8%